OpenStudy (anonymous):

4. Identify the point (x1, y1) from the equation: y + 8 = 3(x – 2) (–2, 8) (8, –2) (–8,2) (2, –8)

sammixboo (sammixboo):

Looks liek the slope-point form (y - y1) = m(x - x1) Maybe that'll help

OpenStudy (triciaal):

the format is y-y1 = m( x - x1)

OpenStudy (anonymous):

Im still confused. I dont understand math at all.

sammixboo (sammixboo):

And remember, two negative signs next to each other turn into an addition sign, so our y1 is going to be negative

sammixboo (sammixboo):

So that marks out choices A and C as our answers

OpenStudy (anonymous):

I think its B??

OpenStudy (triciaal):